OpenStudy (anonymous):

Prove that \(\forall n \in \mathbb{N}, n \geq 1\) \( \Large \prod_{k=1}^n (2k-1) = \frac{(2n)!}{2^{n}.n!}\) I need just solution for n->n+1

OpenStudy (zzr0ck3r):

so ∏ n k=1 (2k−1)=(2n)! 2 n .n! for some n in N

OpenStudy (zzr0ck3r):

so we have (2(n+1)-1)*((2n)!)/(2^n*n!) <-- you understnad how i got here for n+1? its our inductive hypothisis and the fact that ∏ from k=1 to n+1 is the same thing as (2(n+1)-1)*∏ from k = 1 to n

OpenStudy (anonymous):

yes i understand i done same thing as you, what would be the next step ?

OpenStudy (zzr0ck3r):

(2(n+1)-1)*((2n)!)/(2^n*n!) = (2n+1)(2n)!/(2^n*n!) = (2n+2)(2n+1)(2n!)/(2^n*(2n+2)*n!) = (2n+2)!/(2^n*2*(n+1)*(n!)) = (2(n+1))!/(2^(n+1)*(n+1)!)

OpenStudy (zzr0ck3r):

make sense? I got to go to bed soon

OpenStudy (anonymous):

i have problem after second equal sign

OpenStudy (zzr0ck3r):

just multiplied both top and bottom by (2n+2)

OpenStudy (zzr0ck3r):

I did not just know to do that, I worked the RHS on paper and then showed you the backwards steps to get there.

OpenStudy (zzr0ck3r):

its ok to do this step because 2n+2 will only be 0 when n = -1 and n is in N so this will never happen

OpenStudy (anonymous):

hmm i think its right solution but i need some other way, i had seen the solution of this question by some friend and the end solution was shorter then this one and clearer for me, but thanks for your effort

OpenStudy (zzr0ck3r):

if you stay on the lhs you can get to (2n+1)(2n)!/(2^n*n!) and if you start on the RHS you can get to (2n+2)!/(2^n*2*(n+1)*(n!)) and then (2n+1)(2n)!/(2^n*n!) try to do this on paper and see how it goes from there.

OpenStudy (zzr0ck3r):

good luck man, I think I could write this so its nicer but it would take me an hour with the editor:)

OpenStudy (anonymous):

no problem thank you zzrOrck3r :)

OpenStudy (zzr0ck3r):

hope that helps...

OpenStudy (anonymous):

thanks that looks much better ;)

OpenStudy (zzr0ck3r):

lol:) aight gnight

OpenStudy (anonymous):

thank you if you didnt slept already i have one question :) how this possible? \[ \Large 2^{n+1}.(n+1)! = 2^{n}.2.(n+1)(n!)\]

OpenStudy (anonymous):

for my logic in right side \[ \Large (n!)\] is too much

OpenStudy (zzr0ck3r):

2^(n+1) = 2^n * 2 you cool with that part?

OpenStudy (anonymous):

yes

OpenStudy (zzr0ck3r):

(n+1)! = (n+1)*(n+1-1)! = (n+1)*(n)! just like 3! = 3*(3-1)! = 3 * 2!

OpenStudy (zzr0ck3r):

(n+3)! = (n+3)(n+2)(n+1)n!

OpenStudy (anonymous):

ok zzrOck3r i think you are right, my head is too thick to get sometimes such things ;)

OpenStudy (zzr0ck3r):

its cool we all do that the first time we encounter that issue: but yes its the same thing as 5! = 5*4*3! aight you will be fine:)

OpenStudy (anonymous):

thx :)

OpenStudy (zzr0ck3r):

np

OpenStudy (anonymous):

Here is a proof without induction \[ (2n)! = 1\, 2\, 3\, 4\, 5 \cdots (2n -1)\, 2n= \\ 1\, 3\, , 5 \cdots (2n -1)\, 2\, 4\, \cdots \, 2n= \\ \prod_{k=1}^{n} (2k-1)\, (2)\,1\, (2)\,2\, (2)\,3 \cdots \,(2)n=\\ \prod_{k=1}^{n} (2k-1)\, 2^n n!\\ \\ \frac{(2n)!}{2^n n!}=\prod_{k=1}^{n} (2k-1) \]

OpenStudy (anonymous):

thank you mr Saab, your answer helped

OpenStudy (anonymous):

yw